show that the area of parallelogram formed by the lines 2x-3y+a=0,3x-2y-a=0,2x-3y+3a=0 and 3x-2y-2a=0 is 2a²/5 square units

Grade:11

2 Answers

Elton
6 years ago
there is a direct frmula to solve this problem.
Area of paralellogram formed by the lines a1x+b1y+c1=0,a2x+b2y+d1=0,a1x+b1y+c2=0 and a2x+b2y+d2=0 is given by:
Area=(c1-c2)(d1-d2)/(a1b2-a2b1)
Pushpam Kumar PK
6 years ago
Can you solve.... I`m not getting... Please.....,,solve and explain it whole......I wanna pure explanation......